30/x = 9/5
What is the value of x?

Answer by cwmcwilly07(5)

Just simply finding the ratio from one to another
30:9 = 30/9 = 3.33333
Now take that number times the other value
5*3.33333 = 16.66667
30/16.66667 = 9/5